OpenStudy (anonymous):

What are the vertical asymptotes of 1/cos x?

OpenStudy (goformit100):

@ParthKohli

OpenStudy (goformit100):

sec x

OpenStudy (anonymous):

Whenever \(\cos(x)=0\). Try solving that.

Parth (parthkohli):

\[\cos (x) = 0 \implies x = {\pi \over 2} , {3\pi \over 2}, {5\pi \over 2}\cdots\]